# TextureSplat: Per-Primitive Texture Mapping for Reflective Gaussian Splatting

### v1

@@ 4. Experiments
| MAE∘ | 22.31 | 17.02 | 4.618 | 4.871 | 2.078 | 1.83 |
Following our baseline method (Yao et al., 2025), we evaluate our work quantitatively and qualitatively under standard multi-view reconstruction benchmarks of challenging reflective scenes. We use the datasets: Shiny Blender (Verbin et al., 2022) and Glossy Synthetic (Liu et al., 2023) for novel view synthesis of reflective objects, and dataset Ref-Real (Verbin et al., 2022) to account for real world open reflective scenes. We also provide results on the Synthetic NeRF (Mildenhall et al., 2020) dataset to showcase our method under non-reflective scenes and illustrate its practicality. We compare to state-of-the-art methods in the reflective scene setting, including method Ref-Gaussian (Yao et al., 2025), with whom we share the same underlying physics based model and training strategy, competitive reflective Gaussian splatting based methods GShader (Jiang et al., 2024a) and 3DGS-DR (Ye et al., 2024), 3DGS (Kerbl et al., 2023) and 2DGS (Huang et al., 2024) for reference, and seminal NeRF based approaches such as Ref-NeRF (Verbin et al., 2022) and ENVIDR (Liang et al., 2023). We provide additional results and ablation studies in the supplementary material.
4.1. Implementation Details
For a stable optimization, we follow a two-stage optimization approach, where we first train for half the total number of iterations using per-splat single attribute optimization. During the second stage, we start optimizing the material and normal textures initialized from the corresponding attributes from the first stage and we freeze the positions of the primitives.
In all our evaluations, we use texture resolution of 4x4.
We use the same hyper-parameters and training strategies defined by Ref-Gaussian (Yao et al., 2025) as we build on it and consider it our main baseline method.
We also follow Ref-Gaussian (Yao et al., 2025) in replacing the integrated diffuse lighting by a spherical harmonics view dependent color for better fitting in the reflective setting.
We implement efficient CUDA kernels on top of 2DGS and Ref-Gaussian for forward and backward operations involving material textures and normal mapping, as well as the texture atlases construction, packing and hardware bilinear filtering at test time.
4.2. Novel View Synthesis
Table 1 shows numerical results in the standard reflective benchmark. We report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) (Wang et al., 2004), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) as metrics. Our method outperforms the competition, including baseline Ref-Gaussian, which is also the state-of-the-art method currently under this benchmark to the best of our knowledge. We provide qualitative results to accompany this table, for real scenes in Figure 2 and synthetic ones in Figure 3. Our method displays superior ability in capturing reflections on the surface with more fidelity, while competition Ref-Gaussian and 3DGS-DR suffer from distorted or missing reflections. We also show in the supplementary material numerical and qualitative comparisons under non-reflective data on Nerf Synthetic Scenes (Mildenhall et al., 2020). Here our method performs competitively with the competition. The improvement brought by our method over the baseline Ref-Gaussian particularly, under both reflective and non-reflective scenes, is a testimony of the efficacy and versatility of our representation.
4.3. Scene Decomposition
Despite the under-constrained nature of the task in many cases, our optimization yields stable and physically plausible scene property factorization. Figure 5 illustrates decomposition examples for reflective objects into geometry normals, roughness and metallic. We also show the diffuse and specular components, in addition to the final rendering which preserves the groundtruth. The specular highlights are mostly successfully factored out from the view-independent component. Figure 6 shows a comparison of the decomposition with respect to our baseline method. Notice that our material properties are sharper and display less noise. Our normals replicate the smooth sphere shape more faithfully. We compare estimated environment maps for our method and the competition in Figure 4. Notice that our estimated scene lighting bares more similarity with the reference maps and display less distortions and artefacts generally, compared to our baseline.
We evaluate our normal estimation through the benchmark of the Shiny Blender dataset (Verbin et al., 2022). Table 2 reports the mean angular error of normal maps, where we outperform our baseline and the competition. This results validates our tangential normal representation and the use of normal mapping, as also confirmed by our superior qualitative results (Figure 7).
4.4. Hardware Acceleration Performance
@@
We trained our model on the Shiny Blender dataset using three different texture resolutions: 4×4, 8×8, and 16×16. For each model, we then rendered the scenes using both the software implementation of bilinear filtering and our hardware-accelerated implementation with texture atlases. We measured the rendering performance in frames per second (FPS) and compared these values to the baseline method without textures.
| 4×4 Textures | 8×8 Textures | 16×16 Textures | |
| Baseline (No Textures) | 1.00× | 1.00× | 1.00× |
| Software Bilinear | 0.95× | 0.96× | 0.85× |
| Hardware Texture Atlas | 0.96× | 0.97× | 0.93× |
As shown in Table 3, software-based bilinear filtering introduces some performance overhead compared to the baseline. This overhead increases with higher texture resolutions.
In contrast, our hardware-accelerated implementation with texture atlases maintains rendering performance very close to the baseline method in comparison. The performance remains strong even at higher texture resolutions.
These results demonstrate that the texture atlas approach effectively leverages GPU hardware capabilities to minimize the performance impact of texture filtering. Several factors contribute to this efficiency including the use of dedicated texture units that are designed specifically for texture filtering operations, and are substantially faster than general-purpose compute. The gap can be widened even more when applied to more complicated objects or larger scenes.
@@
By leveraging hardware acceleration through texture atlases, we can achieve the best of both worlds: the improved rendering quality of per-primitive textures for reflective scenes while maintaining rendering performance comparable to methods without textures.
4.5. Ablation Study
To evaluate the effectiveness of our per-primitive texture mapping approach and understand the contribution of individual components, we conducted a series of ablation studies. These experiments isolate different aspects of our method and compare against carefully designed baselines to provide insights into why our approach works well for reflective scenes. More experiments can be found in the supplementary material.
@@ 4.5. Ablation Study
| PSNR | 34.89 | 35.47 | 35.94 |
| SSIM | 0.974 | 0.975 | 0.977 |
| LPIPS | 0.053 | 0.051 | 0.047 |
| Normal MAE | 2.078 | 1.845 | 1.825 |
@@
While normal mapping provides the most significant improvement, our full model with texture mapping for all material properties achieves the best overall performance, demonstrating that each component contributes to the final quality.
5. Limitations and Discussion
While our per-primitive texture mapping approach significantly enhances the representation power of 2D Gaussian Splatting for the reflective setting, it comes with several limitations that present opportunities for future work:
Storage Overhead
Storing material textures for each primitive introduces additional memory requirements. Although we save some memory by encoding tangent normals with only two components, the overall storage increases with the number of primitives and texture resolution.
Dependency on Underlying Rendering Model
Our method is designed to be generic and applicable to any Gaussian splatting variant with material attributes. While it enhances reconstruction quality through increased expressivity, it cannot overcome fundamental limitations of the underlying rendering model. For scenes where the physical assumptions of the model are violated (e.g., complex shadowing, grazing angle reflections, multiple level of reflections), our approach may struggle to produce accurate results. In some cases, the additional degrees of freedom provided by normal mapping can lead to overfitting to noise when the underlying model cannot explain the observed data. This explains why we can observe more substantial improvements for simpler reflective scenes compared to more complex ones sometimes.
@@ 5. Limitations and Discussion
In the current implementation, we assign textures of uniform resolution to all primitives regardless of their size or importance in the scene. This approach is inefficient for large unbounded scenes where distant primitives occupy few pixels but still receive the same texture resolution as foreground elements. A more sophisticated approach would involve adaptive texture allocation based on primitive size, viewing distance, or local detail complexity. Primitives in the background could potentially use single attribute values while reserving detailed textures for foreground elements.
@@
Our implementation currently relies on bilinear filtering for texture sampling. While this works well for the scenes in our evaluation, it does not fully address texture minification artifacts that might appear in more challenging scenarios. Implementing proper mipmapping with trilinear filtering would be beneficial but challenging to integrate with the non-antialiased nature of 2D Gaussian primitives. Potential solutions could involve combining our approach with screen-space antialiasing techniques or developing specialized filtering methods for Gaussian-based representations.
Despite these limitations, our experimental results demonstrate that per-primitive texture mapping significantly improves the visual quality of reflective scenes while maintaining real-time rendering performance. The ability to capture fine normal details through normal mapping proves particularly valuable for specular reflections, where small normal variations can have substantial visual impact. Furthermore, by leveraging hardware-accelerated texture filtering through our atlas-based implementation, we achieve these quality improvements with minimal performance overhead compared to traditional Gaussian splatting approaches.
6. Conclusion
@@
We presented a method that enhances 2D Gaussian Splatting for reflective scenes by introducing per-primitive texture mapping. By leveraging the flat nature of 2D Gaussians to define textures of material properties, our approach enables high-frequency detail representation without increasing primitive count. Our hardware-accelerated implementation using texture atlases demonstrates that classical computer graphics techniques can be effectively integrated with modern differentiable rendering approaches.
The results show that this representation significantly improves the quality of specular reflections, particularly through detailed normal mapping, while maintaining real-time performance. Our work bridges the gap between explicit primitive-based representations and high-quality material modeling, offering advantages of both approaches.
Future work could explore several promising directions like advanced filtering techniques like anisotropic filtering for improved quality at grazing angles and compressed texture formats to reduce memory requirements.
Acknowledgements.
@@
This work was granted access to the HPC resources of IDRIS under the allocation 20XX-AD010616156 made by GENCI.References
- (1)
- Burley and Lacewell (2008) Brent Burley and Dylan Lacewell. 2008. Ptex: Per-Face Texture Mapping for Production Rendering. In Proceedings of Eurographics Symposium on Rendering. Eurographics Association. https://graphics.pixar.com/library/Ptex/paper.pdf
- Chao et al. (2025) Brian Chao, Hung-Yu Tseng, Lorenzo Porzi, Chen Gao, Tuotuo Li, Qinbo Li, Ayush Saraf, Jia-Bin Huang, Johannes Kopf, Gordon Wetzstein, and Changil Kim. 2025. Textured Gaussians for Enhanced 3D Scene Appearance Modeling. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). https://arxiv.org/abs/2411.18625
- Chen et al. (2022) Anpei Chen, Zexiang Xu, Andreas Geiger, Jingyi Yu, and Hao Su. 2022. TensoRF: Tensorial Radiance Fields. In Proceedings of the European Conference on Computer Vision (ECCV).
- Fridovich-Keil et al. (2023) Sara Fridovich-Keil, Giacomo Meanti, Frederik Rahbæk Warburg, Benjamin Recht, and Angjoo Kanazawa. 2023. K-Planes: Explicit Radiance Fields in Space, Time, and Appearance. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR).
- Fridovich-Keil et al. (2022) Sara Fridovich-Keil, Alex Yu, Matthew Tancik, Qinhong Chen, Benjamin Recht, and Angjoo Kanazawa. 2022. Plenoxels: Radiance fields without neural networks. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 5501–5510.
- Gao et al. (2024) Jian Gao, Chun Gu, Youtian Lin, Zhihao Li, Hao Zhu, Xun Cao, Li Zhang, and Yao Yao. 2024. Relightable 3D Gaussians: Realistic Point Cloud Relighting with BRDF Decomposition and Ray Tracing. In Proceedings of the European Conference on Computer Vision (ECCV). https://arxiv.org/abs/2311.16043
- Ge et al. (2023) Wenhang Ge, Tao Hu, Haoyu Zhao, Shu Liu, and Ying-Cong Chen. 2023. Ref-NeuS: Ambiguity-Reduced Neural Implicit Surface Learning for Multi-View Reconstruction with Reflection. In Proceedings of the IEEE/CVF International Conference on Computer Vision (ICCV). IEEE, 4251–4260.
- Gouraud (1971) Henri Gouraud. 1971. Continuous shading of curved surfaces. In Proceedings of the 2nd conference on Computer graphics and interactive techniques. ACM, 747–755.
- Huang et al. (2024) Binbin Huang, Zehao Yu, Anpei Chen, Andreas Geiger, and Shenghua Gao. 2024. 2d gaussian splatting for geometrically accurate radiance fields. In ACM SIGGRAPH 2024 conference papers. 1–11.
- Huang and Gong (2024) Zhentao Huang and Minglun Gong. 2024. Textured-GS: Gaussian Splatting with Spatially Defined Color and Opacity. arXiv preprint arXiv:2407.09733 (2024). https://arxiv.org/abs/2407.09733
- Jiang et al. (2024a) Yingwenqi Jiang, Jiadong Tu, Yuan Liu, Xifeng Gao, Xiaoxiao Long, Wenping Wang, and Yuexin Ma. 2024a. GaussianShader: 3D Gaussian Splatting with Shading Functions for Reflective Surfaces. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, 1185–1194. https://doi.org/10.1109/CVPR.2024.00127
- Jiang et al. (2024b) Zijie Jiang, Tianhan Xu, and Hiroharu Kato. 2024b. Rethinking Directional Parameterization in Neural Implicit Surface Reconstruction. In Proceedings of the European Conference on Computer Vision (ECCV). Springer, 127–142.
- Kerbl et al. (2023) Bernhard Kerbl, Georgios Kopanas, Thomas Leimkühler, and George Drettakis. 2023. 3d gaussian splatting for real-time radiance field rendering. ACM Trans. Graph. 42, 4 (2023), 139–1.
- Kingma and Ba (2015) D.P. Kingma and J.B. Ba. 2015. Adam: A Method for Stochastic Optimization. International Conference on Learning Representations (ICLR) (2015). https://arxiv.org/abs/1412.6980
- Levoy (1988) Marc Levoy. 1988. Display of Surfaces from Volume Data. IEEE Computer Graphics and Applications 8, 3 (1988), 29–37.
- Li et al. (2023) Zhaoshuo Li, Thomas Müller, Alex Evans, Russell H Taylor, Mathias Unberath, Ming-Yu Liu, and Chen-Hsuan Lin. 2023. Neuralangelo: High-fidelity neural surface reconstruction. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 8456–8465.
- Liang et al. (2023) Ruofan Liang, Huiting Chen, Chunlin Li, Fan Chen, Selvakumar Panneer, and Nandita Vijaykumar. 2023. ENVIDR: Implicit Differentiable Renderer with Neural Environment Lighting. In Proceedings of the IEEE/CVF International Conference on Computer Vision (ICCV). IEEE, 79–89.
- Liang et al. (2024) Zhihao Liang, Qi Zhang, Ying Feng, Ying Shan, and Kui Jia. 2024. GS-IR: 3D Gaussian Splatting for Inverse Rendering. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). 21644–21653.
- Liu et al. (2020) Lingjie Liu, Jiatao Gu, Kyaw Zaw Lin, Tat-Seng Chua, and Christian Theobalt. 2020. Neural Sparse Voxel Fields. In NeurIPS.
- Liu et al. (2023) Yuan Liu, Peng Wang, Cheng Lin, Xiaoxiao Long, Jiepeng Wang, Lingjie Liu, Taku Komura, and Wenping Wang. 2023. NeRO: Neural Geometry and BRDF Reconstruction of Reflective Objects from Multiview Images. ACM Trans. Graph. 42, 4, Article 114 (jul 2023), 22 pages.
- Max (1995) Nelson Max. 1995. Optical models for direct volume rendering. IEEE Transactions on Visualization and Computer Graphics 1, 2 (1995), 99–108.
- Mildenhall et al. (2020) Ben Mildenhall, Pratul P. Srinivasan, Matthew Tancik, Jonathan T. Barron, Ravi Ramamoorthi, and Ren Ng. 2020. NeRF: Representing Scenes as Neural Radiance Fields for View Synthesis. In ECCV. Springer, 405–421.
- Müller et al. (2022) Thomas Müller, Alex Evans, Christoph Schied, and Alexander Keller. 2022. Instant neural graphics primitives with a multiresolution hash encoding. ACM transactions on graphics (TOG) 41, 4 (2022), 1–15.
- Phong (1975) Bui Tuong Phong. 1975. Illumination for computer generated pictures. Commun. ACM 18, 6 (1975), 311–317.
- Purnomo et al. (2004) Budirijanto Purnomo, Jonathan D Cohen, and Subodh Kumar. 2004. Seamless texture atlases. In IEEE Visualization, 2004. VIS 2004. IEEE, 273–280. https://doi.org/10.1109/VISUAL.2004.91
- Srinivasan et al. (2021) Pratul P. Srinivasan, Boyang Deng, Xiuming Zhang, Matthew Tancik, Ben Mildenhall, and Jonathan T. Barron. 2021. NeRV: Neural Reflectance and Visibility Fields for Relighting and View Synthesis. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). 7495–7504. https://doi.org/10.1109/CVPR46437.2021.00741
- Sun et al. (2024) Cheng Sun, Jaesung Choe, Charles Loop, Wei-Chiu Ma, and Yu-Chiang Frank Wang. 2024. Sparse Voxels Rasterization: Real-time High-fidelity Radiance Field Rendering. arXiv preprint arXiv:2412.04459 (2024).
- Sun et al. (2022) Jingxiang Sun, Yiming Gao, Xuan Wang, Qi Zhang, Hujun Bao, and Xiaowei Zhou. 2022. Improved Direct Voxel Grid Optimization for Radiance Fields. arXiv:2206.05085 [cs.CV]
- Tang and Cham (2024) Zhe Jun Tang and Tat-Jen Cham. 2024. 3iGS: Factorised Tensorial Illumination for 3D Gaussian Splatting. In Proceedings of the European Conference on Computer Vision (ECCV). Springer. https://doi.org/10.1007/978-3-031-72630-9_9
- Verbin et al. (2022) Dor Verbin, Peter Hedman, Ben Mildenhall, Todd Zickler, Jonathan T Barron, and Pratul P Srinivasan. 2022. Ref-nerf: Structured view-dependent appearance for neural radiance fields. In Proceedings of the IEEE/CVF CVPR. IEEE, 5481–5490.
- Wang et al. (2021) Peng Wang, Lingjie Liu, Yuan Liu, Christian Theobalt, Taku Komura, and Wenping Wang. 2021. Neus: Learning neural implicit surfaces by volume rendering for multi-view reconstruction. arXiv preprint arXiv:2106.10689 (2021).
- Wang et al. (2004) Zhou Wang, Alan C Bovik, Hamid R Sheikh, and Eero P Simoncelli. 2004. Image quality assessment: from error visibility to structural similarity. IEEE transactions on image processing 13, 4 (2004), 600–612.
- Wu et al. (2022) Tong Wu, Jiaqi Wang, Xingang Pan, Xudong Xu, Christian Theobalt, Ziwei Liu, and Dahua Lin. 2022. Voxurf: Voxel-based efficient and accurate neural surface reconstruction. arXiv preprint arXiv:2208.12697 (2022).
- Xu et al. (2024) Tian-Xing Xu, Wenbo Hu, Yu-Kun Lai, Ying Shan, and Song-Hai Zhang. 2024. Texture-GS: Disentangling the Geometry and Texture for 3D Gaussian Splatting Editing. In Proceedings of the European Conference on Computer Vision (ECCV). Springer. https://arxiv.org/abs/2403.10050
- Yao et al. (2025) Yuxuan Yao, Zixuan Zeng, Chun Gu, Xiatian Zhu, and Li Zhang. 2025. Reflective Gaussian Splatting. In ICLR.
- Yariv et al. (2021) Lior Yariv, Jiatao Gu, Yoni Kasten, and Yaron Lipman. 2021. Volume rendering of neural implicit surfaces. Advances in Neural Information Processing Systems 34 (2021), 4805–4815.
- Ye et al. (2024) Keyang Ye, Qiming Hou, and Kun Zhou. 2024. 3D Gaussian Splatting with Deferred Reflection. In ACM SIGGRAPH Conference Proceedings. https://doi.org/10.1145/3641519.3657456
- Zhang et al. (2018) Richard Zhang, Phillip Isola, Alexei A Efros, Eli Shechtman, and Oliver Wang. 2018. The unreasonable effectiveness of deep features as a perceptual metric. In Proceedings of the IEEE conference on computer vision and pattern recognition. 586–595.
- Zhu et al. (2024) Zuo-Liang Zhu, Beibei Wang, and Jian Yang. 2024. GS-ROR: 3D Gaussian Splatting for Reflective Object Relighting via SDF Priors. arXiv preprint arXiv:2406.18544 (2024). https://arxiv.org/abs/2406.18544
- Zwicker et al. (2001) Matthias Zwicker, Hanspeter Pfister, Jeroen Van Baar, and Markus Gross. 2001. EWA Volume Splatting. In Proceedings of the IEEE Conference on Visualization (VIS). IEEE, 29–36.
- Zwicker et al. (2002) Matthias Zwicker, Hanspeter Pfister, Jeroen Van Baar, and Markus Gross. 2002. EWA splatting. IEEE Transactions on Visualization and Computer Graphics 8, 3 (2002), 223–238.
- Zwicker et al. (2004) Matthias Zwicker, Jussi Rasanen, Mario Botsch, Carsten Dachsbacher, and Mark Pauly. 2004. Perspective accurate splatting. In Proceedings-Graphics Interface. 247–254.
7. Additional Ablation Studies
Texture Mapping vs. Increased Primitive Count
A natural question when evaluating our texture-based method is whether similar quality improvements could be achieved by simply increasing the number of primitives in the baseline method. To investigate this, we compared our approach against baseline models with increased primitive counts. We created versions of the baseline with 2× and 4× the original number of primitives by lowering the gradient threshold for densification and increasing the frequency of cloning and splitting operations during optimization. We compare against our method that maintain the original primitive count with the 4×4 textures per primitive. Notably, even with these textures, our method has lower storage requirements than the 2× primitive count baseline. As shown in Table 5, the performance of the baseline method actually degrades as the number of primitives increases beyond the optimal point. This counter-intuitive result can be explained by the fact that increasing the number of primitives creates a more complex optimization and that more primitives require more extensive sorting operations during rendering, which can lead to primitives being placed in suboptimal orders, especially with the use of global sorting (Kerbl et al., 2023). Primitives can also become cluttered, particularly in areas with complex geometry, leading to degraded normal estimation quality which is critical for reflective scenes.
| Metric | Ours | Ref-Gaussian | 2 Primitives | 4 Primitives |
| PSNR | 35.94 | 34.89 | 33.63 | 34.12 |
| SSIM | 0.977 | 0.974 | 0.967 | 0.970 |
| LPIPS | 0.047 | 0.053 | 0.061 | 0.057 |
Performance with Reduced Primitive Count
As illustrated in Figure 8, the baseline method with reduced primitive count suffers from significant degradation in normal reconstruction quality, exhibiting holes and artifacts due to the limited capacity of the model. In contrast, our texture-based method maintains smooth normal reconstruction even with the reduced primitive count, demonstrating the expressiveness of per-primitive texture mapping.
This experiment highlights a key advantage of our approach: it can achieve high-quality results with fewer primitives, making it particularly suitable for applications with strict performance requirements or memory constraints.

### v2

@@ 4.3 Scene Decomposition
| MAE∘ | 22.31 | 17.02 | 4.618 | 4.871 | 2.078 | 1.78 |
4.4 Hardware Acceleration Performance
@@
We trained our model on the Shiny Blender dataset using different texture resolutions: 2×2, 4×4, 8×8 and 16×16. For each model, we then rendered the scenes using both the software implementation of bilinear filtering and our hardware-accelerated implementation with texture atlases. We measured the rendering performance in frames per second (FPS) and compared these values to the baseline method without textures.
| 2×2 Textures | 4×4 Textures | 8×8 Textures | 16×16 Textures | |
| Baseline, No Textures | 1.00 | 1.00 | 1.00 | 1.00 |
| Software Bilinear | 0.90 | 0.86 | 0.85 | 0.79 |
| Hardware Texture Atlas | 0.92 | 0.93 | 0.94 | 0.91 |
As shown in Table 3, software-based bilinear filtering introduces some performance overhead compared to the baseline. This overhead increases with higher texture resolutions. In contrast, our hardware-accelerated implementation with texture atlases maintains rendering performance very close to the baseline method in comparison.
These results demonstrate that the texture atlas approach effectively leverages GPU hardware capabilities to minimize the performance impact of texture filtering. Several factors contribute to this efficiency including the use of dedicated texture units that are designed specifically for texture filtering operations, and are substantially faster than general-purpose compute.
@@
By leveraging hardware acceleration through texture atlases, we can achieve the best of both worlds: the improved rendering quality of per-primitive textures for reflective scenes while maintaining rendering speed comparable to methods without textures.
4.5 Ablation Study
To evaluate the effectiveness of our per-primitive texture mapping approach and understand the contribution of individual components, we conducted a series of ablation studies isolating different aspects of our method.
Additional experiments can be found in Sec. 3 of the supplementary material, notably a comparative analysis of Texture Mapping as opposed to increasing primitive count, and a comparative analysis of performance under reduced primitive count (Tab. 2 & Fig. 4 in Supp. Mat.).
@@ 4.5 Ablation Study
| PSNR | 34.89 | 35.33 | 35.90 |
| SSIM | 0.974 | 0.975 | 0.978 |
| LPIPS | 0.053 | 0.053 | 0.047 |
| Normal MAE | 2.078 | 1.783 | 1.78 |
| Variants | PSNR | SSIM | LPIPS | Normal MAE | Storage |
| Ref-Gaussian | 34.88 | 0.974 | 0.053 | 2.078 | 1.00 |
| 22 Textures, Same Storage | 35.66 | 0.976 | 0.048 | 1.823 | 1.00 |
| 22 Textures | 35.90 | 0.977 | 0.047 | 1.780 | 1.20 |
| 44 Textures, Same Storage | 35.33 | 0.977 | 0.048 | 1.881 | 1.00 |
| 44 Textures | 35.94 | 0.978 | 0.047 | 1.830 | 1.96 |
@@
While normal mapping provides significant improvement, our full model with texture mapping for all material properties achieves the best overall performance, demonstrating that each component contributes to the final quality.
Impact of Texture Resolution and Storage
The results in Table 5 show different configurations of our method in terms of the texture resolution, and whether we add them on top of the baseline or reduce the primitive count so that the total storage size is equal to the baseline. We perform this evaluation on the Shiny Blender dataset [ref_nerf]. We experiment with resolutions 22 and 44 as we find that for the type of scenes in this dataset (object at the center of the scene and orbital views), enhancing primitives with much higher texture resolutions does not result in a desirable storage to quality tradeoff. Using higher resolutions can still be beneficial though for other scene types where some views are close to the object for instance.
We find that at a fixed storage budget, our method still achieves better performance compared to the baseline method in all metrics. This demonstrates the efficiency of our texture-based representation and that the memory overhead is not a fixed cost but a tunable parameter. Our framework allows for adapting texture resolutions or selectively texturing only the most critical attributes to fit different memory budgets.
Supplementary material includes additional comparisons to the baseline with increased primitive count as well as the impact on performance with reduced primitive count; These experiments show that for reflective scenes, investing the parameter budget in appearance/normal complexity (textures) is often more effective than investing it purely in geometric complexity (primitives).
5 Limitations and Discussion
While our per-primitive texture mapping approach significantly enhances the representation power of 2D Gaussian Splatting for the reflective setting, it comes with some limitations that present opportunities for future work.
@@ 5 Limitations and Discussion
In the current implementation, we assign textures of uniform resolution to all primitives regardless of their size or importance in the scene. This approach can be inefficient for large unbounded scenes where distant primitives occupy few pixels but still receive the same texture resolution as foreground elements. A more sophisticated approach would involve adaptive texture allocation based on primitive size, viewing distance, or local detail complexity. Primitives in the background could potentially use single attribute values while reserving detailed textures for more predominant foreground elements.
@@
Our implementation currently relies on bilinear filtering for texture sampling. While effective for our evaluated scenes, it does not fully resolve texture minification artifacts that might manifest in more challenging scenarios. while implementing proper mipmapping with trilinear filtering could be beneficial, it would be challenging to integrate with the non-antialiased nature of 2D Gaussian primitives. Potential solutions could involve combining our approach with screen-space antialiasing techniques or developing specialized filtering methods for Gaussian-based representations.
Despite these limitations, our experimental results demonstrate that per-primitive texture mapping significantly improves the visual quality of reflective scenes while maintaining real-time rendering performance.
6 Conclusion
@@
We presented a method that enhances 2D Gaussian Splatting for reflective scenes by introducing per-primitive texture mapping. By leveraging the flat nature of 2D Gaussians to define textures of material properties, our approach enables high-frequency detail representation without increasing primitive count. Our hardware-accelerated implementation using texture atlases demonstrates that classical computer graphics techniques can be effectively integrated with modern differentiable rendering approaches. The results show that this representation significantly improves the quality of specular reflections, particularly through detailed normal mapping, while maintaining real-time performance. Our work bridges the gap between explicit primitive-based representations and high-quality material modeling, offering advantages of both approaches.
Acknowledgments
@@
This work was granted access to the HPC resources of IDRIS under the allocation 20XX-AD010616156 made by GENCI.
Supplementary Material
7 Implementation Details of Hardware Accelerated Texture Atlases
Texture Atlas Construction
Each splat is associated with a small, fixed-resolution texture (e.g., texels, where is typically in powers of ). For splats and total channels for all textured attributes, this texel data is organized into one or more large 2D atlas textures. We pack the individual primitive textures into larger texture atlases that can be efficiently accessed by the GPU by determining the maximum atlas dimensions based on GPU capabilities, calculating how many textures fit along each dimension, computing the total number of textures per atlas, then determining the number of atlases needed.
Hardware Accelerated Texture Sampling
1.
For a given primitive and local coordinates , we map to normalized texture coordinates as described in the main paper (Section 3.3).
2.
We retrieve the chart coordinates from the indirection buffer.
3.
We compute the atlas coordinates:
(14) where and are the total number of charts along the width and height of the atlas, respectively.
4.
We sample the atlas using hardware-accelerated bilinear filtering with CUDA’s ‘tex2D()‘ intrinsic: